solve the system:
{x+5y-z=4
{3x+4y+z=4
{x-y+z=0.
thanks for the help.

Answer by stanbon(75887)

{x+5y-z=4
{3x+4y+z=4
{x-y+z=0.
----------------
Using a TI-83 Matrix function:
x = 2
y = 0
z = -2
------------
Cheers,
Stan H.